# Comparing Computational Pathology Foundation Models using Representational Similarity Analysis

### v1

@@ Abstract
Foundation models are increasingly developed in computational pathology (CPath) given their promise in facilitating many downstream tasks. While recent studies have evaluated task performance across models, less is known about the structure and variability of their learned representations. Here, we systematically analyze the representational spaces of six CPath foundation models using techniques popularized in computational neuroscience. The models analyzed span vision-language contrastive learning (CONCH, PLIP, KEEP) and self-distillation (UNI (v2), Virchow (v2), Prov-GigaPath) approaches. Through representational similarity analysis using H&E image patches from TCGA, we find that UNI2 and Virchow2 have the most distinct representational structures, whereas Prov-Gigapath has the highest average similarity across models. Having the same training paradigm (vision-only vs. vision-language) did not guarantee higher representational similarity. The representations of all models showed a high slide-dependence, but relatively low disease-dependence. Stain normalization decreased slide-dependence for all models by a range of 5.5% (CONCH) to 20.5% (PLIP). In terms of intrinsic dimensionality, vision-language models demonstrated relatively compact representations, compared to the more distributed representations of vision-only models. These findings highlight opportunities to improve robustness to slide-specific features, inform model ensembling strategies, and provide insights into how training paradigms shape model representations. Our framework is extendable across medical imaging domains, where probing the internal representations of foundation models can help ensure effective development and deployment.
@@ 1 Introduction
Numerous foundation models have recently been developed for computational pathology. Trained on large-scale datasets in a self-supervised fashion, these models are often designed to generate patch-level embeddings from H&E-stained whole slide images (WSIs), which have proven valuable for downstream tasks such as tumor subtyping, biomarker prediction, and prognostication (Bilal et al., 2025; Li et al., 2025; Neidlinger et al., 2024). However, recent analyses have also raised questions regarding the robustness of these models to data-specific confounders, such as hospital-specific signatures (de Jong et al., 2025; Kömen et al., 2024). Furthermore, while studies have compared downstream performance across models, the extent to which these models produce similar or divergent embeddings remains unclear. Given the emphasis of interpretability and robustness in medical AI, there is a pressing need to move beyond performance metrics and systematically examine the representations learned by foundation models in computational pathology (CPath).
Here, we leverage techniques from computational neuroscience to explore the representational spaces of six CPath foundation models. Using representational similarity analysis (RSA) (Kriegeskorte et al., 2008), we examine how H&E-patch embeddings vary across models, diseases, and individual WSIs. The models analyzed represent the two predominant self-supervised learning strategies in the field: vision-language contrastive learning and vision-only self-distillation. As a comparative baseline, we include a DinoV2 model (Oquab et al., 2024) trained on natural images. We aimed to quantify similarities and differences between these models, including characterizing the slide- and disease-dependence of their representations, as well as their intrinsic dimensionality.
@@ 2.1 Foundation models in computational pathology
Following strategies developed for natural images, the two most predominant paradigms for developing CPath foundation models are vision-language contrastive learning and vision-only self-distillation (Bilal et al., 2025; Li et al., 2025). Vision-language contrastive learning, pioneered by CLIP (Radford et al., 2021), aims to learn joint representations between images and text such that matching image-text pairs have similar embeddings and non-matches have dissimilar embeddings. Vision-language models developed for CPath have used various sources to curate image-text pairs, including Twitter (e.g. PLIP (Huang et al., 2023)), Youtube (e.g., Quilt-1M (Ikezogwo et al., 2023)), and PubMed (e.g., CONCH (Lu et al., 2024)). KEEP extended traditional vision-language contrastive learning by creating a knowledge graph to enable hierarchical semantic alignment (Zhou et al., 2024). For self-distillation, the most commonly adopted algorithm is DinoV2 (Oquab et al., 2024), which combines contrastive data augmentation and masking in a student-teacher framework. This learning paradigm aims to learn representations that capture meaningful features of the underlying data that are stable under transformations. CPath models developed using DinoV2 include UNI (Chen et al., 2024), Prov-GigaPath (Xu et al., 2024), and Virchow (Vorontsov et al., 2024; Zimmermann et al., 2024).
@@ 2.2 Analysis of CPath foundation models
Several studies have evaluated the performance of CPath foundation models on downstream tasks such as disease classification, biomarker prediction, and prognostication (Wölflein et al., 2024; Bilal et al., 2025; Neidlinger et al., 2024; Bareja et al., 2025; Marza et al., 2025). In terms of representation analysis, Kömen et al. (2024) demonstrated that the originating data site can be reliably predicted from the embeddings of several vision-only CPath foundation models, a finding also supported by de Jong et al. (2025).
@@ 2.3 Representational similarity analysis
Representational similarity analysis (RSA) is a framework often used in computational neuroscience to compare neural representations across different conditions, such as comparing the responses of neurons to different stimuli (Kriegeskorte et al., 2008). The core of RSA involves constructing representational dissimilarity matrices (RDMs), which capture pairwise distances between representations across conditions. The RDMs for different systems (e.g., brain regions) can then also be compared to offer insights into similarities and differences between representational structures.
@@ 3.1 Models Studied
We selected six popular foundation models for analysis (Table 1), consisting of three vision-only self-distillation models (UNI (v2), Virchow (v2), Prov-Gigapath) and three vision-language models (CONCH, PLIP, KEEP). We additionally include a Dinov2 model trained on natural images as a baseline (Oquab et al., 2024). UNI was trained on over 200 million image patches from over 350k slides sourced from Mass General Brigham (Chen et al., 2024). Virchow was trained on 3.1 million WSIs from the Memorial Sloan Kettering Cancer Center (Zimmermann et al., 2024). For both UNI and Virchow, we use the version 2 models. Prov-GigaPath was trained on 1.3 billion image patches from 171,189 WSIs from the Providence health system (Xu et al., 2024). CONCH was trained on over 1.17 million image–caption pairs curated from various sources, including PubMed (Lu et al., 2024). PLIP was trained on 208,414 image–caption pairs curated from public sources (OpenPath), including Twitter (Huang et al., 2023). KEEP (Zhou et al., 2024) was trained using OpenPath and Quilt-1M (Ikezogwo et al., 2023), a dataset of 1M image–caption pairs curated from public sources, including Youtube. All of the studied models use a form of vision transformer (ViT) for the embedder architecture (Table 1).
@@
UNI Vision ViT-H Mass General Brigham
Virchow ViT-H MSKCC
@@ 3.2 Dataset
We leverage four subsets of TCGA to perform the analysis: breast invasive carcinoma (BRCA), lung adenocarcinoma (LUAD), lung squamous cell carcinoma (LUSC), and colon adenocarcinoma (COAD) (Cancer Genome Atlas Research Network et al., 2013). TCGA was chosen given its common use for benchmarking in computational pathology and because it was not used for pre-training for any of the studied models. Within each cancer subset, 250 whole slide images (WSIs) were randomly sampled to be included in the analysis. For the representational similarity analysis, these WSIs were split into 5 non-overlapping batches of 50 WSIs each. For each WSI, 50 image patches (224x224 pixels) were then sampled following foreground detection using Otsu’s algorithm (Otsu, 1979). Otsu’s method was applied to a grayscale version of the WSI, downsampled by a factor of 224, to identify foreground regions for the patch selection. Altogether the similarity analysis was performed on 50,000 images (4 cancer types * 250 WSIs * 50 patches/WSI), split into 5 batches of 10,000 images each (50 WSIs/batch). The sample size and batching strategy was chosen because of the high computational costs of RSA, which grows according to $N^{2}$, and to provide a measure of uncertainly as quantified by the range across batches.
@@ 3.4 Representational Dissimilarity Matrix Generation
Representational dissimilarity matrix (RDM) construction consists of computing pairwise distances between groups of embeddings. We computed an RDM separately for each model using the Euclidean distance as the distance metric between pairs of embeddings. The Euclidean distance was used as it is recommended over correlation-based distances when computing RDMs given increased robustness to noise at the representation level (Botero and Kriegeskorte, 2024; van den Bosch et al., 2025). When plotting the RDMs, the rows and columns were ordered by WSI and disease type (i.e., all patches from a given WSI are grouped together). RDMs were computed separately for each of the 5 data batches. The rsatoolbox Python package (Nili et al., 2014) was used to compute the RDMs.
@@ 3.5 Comparing RDMs across Models
To quantify the similarity between model RDMs, we computed pairwise comparisons across all seven models. For each model pair, we loaded the upper-triangular values (excluding the diagonal) of their RDMs and computed the Spearman rank correlation. This process was performed separately for each of the 5 batches, yielding five similarity values per each model pair. The results are reported as the mean and range across these values. As an aggregate measure for each model, we also computed the average similarity to all other models. We note that the Spearman correlation is a common metric for comparing RDMs, preferred over Pearson correlation because it does not assume linearity (van den Bosch et al., 2025).
To quantitatively group different models according to similarity, we performed hierarchical clustering on the RDM similarity matrices. Ward’s method (Ward Jr., 1963) was used to produce a linkage tree, visualized as a dendrogram.
@@ 3.6 Assessing Disease and Slide Specificity
In more detail, for slide specificity, we computed the distribution of intra-slide distances as pairwise Euclidean distances between all 50 patch embeddings for a given slide, and aggregating across all diseases and slides. Inter-slide distances were computed by measuring Euclidean distances between all combinations of patches between different slides. Cliff’s Delta was computed between the intra- and inter-slide distances, and this process was performed separately for each of the 5 batches. For each model, we reported the mean and range of Cliff’s Delta values as an indicator of slide specificity.
@@ 4.1 Representational Similarity
By converting the representations for each model in a standardized format that is agnostic to feature size, the RDMs enable comparisons of representations across models. Figure 2 summarizes the degree of similarity between each pair of models, quantified as the Spearman correlation between the RDMs of each model. Higher similarities indicate that image patches with embeddings that are farther apart in one model also tend to be farther apart in the representational space of the other model, indicating common representational structures. The highest overall similarity is observed between UNI2 and Prov-Gigapath, with a Spearman correlation of 0.555 between their RDMs (range of 0.531-0.575 across 5 independent batches of $10K$ image patches). This finding is intuitive as UNI2 and Prov-Gigapath are both vision-only models trained using the Dinov2 algorithm. However, not all patterns align with similarities between self-supervised learning strategy. For instance, out of all of the foundation models, Virchow2’s lowest similarity was observed for UNI2, despite both using DinoV2 for training. On average, Virchow2 and UNI2 were the most distinct foundation models in their representational structure, with an mean correlation with the other CPath models of 0.419 and 0.421, respectively (Table 2). Prov-GigaPath had the highest average similarity (0.523). The baseline ViT-Dinov2 trained on natural images had low similarity ($\leq$0.3) to each model except PLIP (0.503). Hierarchical clustering on the Spearman similarity matrix revealed similar patterns (Figure 3), with ViT-Dinov2 and PLIP separating from the other foundation models.
@@ 4.2 Disease- and Slide- Specificity
In addition to model similarity, we evaluated the degree of slide- and disease-specificity for each model by comparing the relative distances between embeddings from the same slide and different slides, and separately from the same disease and different diseases. Cliff’s Delta, a non-parametric measure of the magnitude of difference between two groups, was used to quantify these effects. Ranging from 0 to 1, a Cliff’s Delta greater than 0.43 is regarded as a large effect, with values of 0.28 and 0.11 indicating medium and weak effects, respectively (Vargha and Delaney, 2000).
@@
The results of the slide- and disease-specificity analyses are contained in Table 3. All of the foundation models demonstrate relatively large slide-specificity, indicating that embeddings from patches from the same WSI tend to be closer together than embeddings from different WSIs. The highest level of slide specificity was observed for Prov-Gigapath, with a Cliff’s Delta of 0.762 (range: 0.658-0.807 across the 5 batches), followed closely by UNI2 (0.751, range: 0.685, 0.780). CONCH demonstrated the lowest value of 0.596 (range: 0.501-0.636). All of the foundation models had a higher slide-specificity than the ViT-Dinov2 baseline (0.367). For disease-specificity, the associations were low for each of the models, with a range of 0.120 (Virchow) to 0.210 (KEEP). Thus, while embeddings from the same disease tended to be slightly closer together than embeddings from different diseases, the representations had a much stronger dependency on individual slides. No clear trends were observed regarding vision-language versus vision-only models for either slide- or disease-specificity.

### v2

@@ Abstract
Foundation models are increasingly developed in computational pathology (CPath) given their promise in facilitating many downstream tasks. While recent studies have evaluated task performance across models, less is known about the structure and variability of their learned representations. Here, we systematically analyze the representational spaces of six CPath foundation models using techniques popularized in computational neuroscience. The models analyzed span vision-language contrastive learning (CONCH, PLIP, KEEP) and self-distillation (UNI (v2), Virchow (v2), Prov-GigaPath) approaches. Through representational similarity analysis using H&E image patches from TCGA, we find that UNI2 and Virchow2 have the most distinct representational structures, whereas Prov-Gigapath has the highest average similarity across models. Having the same training paradigm (vision-only vs. vision-language) did not guarantee higher representational similarity. The representations of all models showed a high slide-dependence, but relatively low disease-dependence. Stain normalization decreased slide-dependence for all models by a range of 5.5% (CONCH) to 20.5% (PLIP). In terms of intrinsic dimensionality, vision-language models demonstrated relatively compact representations, compared to the more distributed representations of vision-only models. These findings highlight opportunities to improve robustness to slide-specific features, inform model ensembling strategies, and provide insights into how training paradigms shape model representations. Our framework is extendable across medical imaging domains, where probing the internal representations of foundation models can support their effective development and deployment.
keywords:
computational pathology, foundation models, robustness, representation analysis
Data and Code Availability
This study leverages the publicly available TCGA dataset. Analysis code is available at https://github.com/lotterlab/cpath_rsa.
Institutional Review Board (IRB)
This research does not require IRB approval.
@@ 1 Introduction
Numerous foundation models have recently been developed for computational pathology. Trained on large-scale datasets in a self-supervised fashion, these models are often designed to generate patch-level embeddings from H&E-stained whole slide images (WSIs), facilitating downstream tasks such as tumor subtyping, biomarker prediction, and prognostication (bilal2025review; li2025review; neidlinger2024). However, recent analyses have also raised questions regarding the robustness of these models to data-specific confounders, such as hospital-specific signatures (dejong2025; komen2024). Furthermore, while studies have compared downstream performance across models, the extent to which these models produce similar or divergent embeddings remains unclear. Given the emphasis of interpretability and robustness in medical AI, there is a pressing need to move beyond performance metrics and systematically examine the representations learned by foundation models in computational pathology (CPath).
Here, we leverage techniques from computational neuroscience to explore the representational spaces of six CPath foundation models. Using representational similarity analysis (RSA) (neuro), we examine how H&E-patch embeddings vary across models, diseases, and individual WSIs. The models analyzed represent the two predominant self-supervised learning strategies in the field: vision-language contrastive learning and vision-only self-distillation. As a comparative baseline, we include a DinoV2 model (oquab2024dinov2learningrobustvisual) trained on natural images (i.e., typical, non-medical RGB images such as those in ImageNet). We aimed to quantify similarities and differences between these models, including characterizing the slide- and disease-dependence of their representations, as well as their intrinsic dimensionality. Given the rapidly evolving field, assessing embedding similarity can offers insights into how different training strategies affect the representations learned and guide practical efforts to enhance downstream performance, such as model ensembling and confounder mitigation.
@@ 2.1 Foundation Models in Computational Pathology
Following strategies developed for natural images, the two most predominant paradigms for developing CPath foundation models are vision-language contrastive learning and vision-only self-distillation (bilal2025review; li2025review). Vision-language contrastive learning, pioneered by CLIP (clip), aims to learn joint representations between images and text such that matching image-text pairs have similar embeddings and non-matches have dissimilar embeddings. Vision-language models developed for CPath have used various sources to curate image-text pairs, including Twitter (e.g. PLIP (huang2023visual)), Youtube (e.g., Quilt-1M (ikezogwo2023quilt1m)), and PubMed (e.g., CONCH (lu2024avisionlanguage)). KEEP extended traditional vision-language contrastive learning by creating a knowledge graph to enable hierarchical semantic alignment (keep). For self-distillation, the most commonly adopted algorithm is DinoV2 (oquab2024dinov2learningrobustvisual), which combines contrastive data augmentation and masking in a student-teacher framework. This learning paradigm aims to learn representations that capture meaningful features of the underlying data that are stable under transformations. CPath models developed using DinoV2 include UNI (chen2024uni), Prov-GigaPath (xu2024gigapath), and Virchow (Vorontsov2024; zimmermann2024virchow2).
@@ 2.2 Analysis of CPath Foundation Models
Several studies have evaluated the performance of CPath foundation models on downstream tasks such as disease classification, biomarker prediction, and prognostication (wolflein2024; bilal2025review; neidlinger2024; Bareja2025; marza2025). In terms of representation analysis, komen2024 demonstrated that the originating data site can be reliably predicted from the embeddings of several vision-only CPath foundation models, a finding also supported by dejong2025.
@@ 2.3 Representational Similarity Analysis
Representational similarity analysis (RSA) is a framework used in computational neuroscience to compare neural representations across different conditions, such as comparing the responses of neurons to different stimuli (neuro). The core of RSA involves constructing representational dissimilarity matrices (RDMs), which capture pairwise distances between representations across conditions. The RDMs for different systems (e.g., brain regions) can then also be compared to offer insights into similarities and differences between representational structures.
@@ 3.1 Models Studied
We selected six popular foundation models for analysis (Table 1), consisting of three vision-only self-distillation models (UNI (v2), Virchow (v2), Prov-Gigapath) and three vision-language models (CONCH, PLIP, KEEP). We additionally include a Dinov2 model trained on natural images as a baseline (oquab2024dinov2learningrobustvisual). UNI was trained on over 200 million image patches from over 350k slides sourced from Mass General Brigham (chen2024uni). Virchow was trained on 3.1 million WSIs from the Memorial Sloan Kettering Cancer Center (zimmermann2024virchow2). For both UNI and Virchow, we use the version 2 models. Prov-GigaPath was trained on 1.3 billion image patches from 171,189 WSIs from the Providence health system (xu2024gigapath). CONCH was trained on over 1.17 million image–caption pairs curated from various sources, including PubMed (lu2024avisionlanguage). PLIP was trained on 208,414 image–caption pairs curated from public sources (OpenPath), including Twitter (huang2023visual). KEEP (keep) was trained using OpenPath and Quilt-1M (ikezogwo2023quilt1m), a dataset of 1M image–caption pairs curated from public sources, including Youtube. All of the studied models use a form of vision transformer (ViT) for the embedder architecture (Table 1).
@@
UNI2 Vision ViT-H Mass General Brigham
Virchow2 ViT-H MSKCC
@@ 3.2 Dataset
We leverage four subsets of TCGA to perform the analysis: breast invasive carcinoma (BRCA), lung adenocarcinoma (LUAD), lung squamous cell carcinoma (LUSC), and colon adenocarcinoma (COAD) (tcga). TCGA was chosen given its common use for benchmarking in computational pathology and because it was not used for pre-training in any of the studied models. Within each cancer subset, 250 whole slide images (WSIs) were randomly sampled to be included in the analysis. For the representational similarity analysis, these WSIs were split into 5 non-overlapping batches of 50 WSIs each. For each WSI, 50 image patches (224x224 pixels) were then sampled following foreground detection using Otsu’s algorithm (otsu). Otsu’s method was applied to a grayscale version of the WSI, downsampled by a factor of 224, to identify foreground regions for the patch selection. Altogether the similarity analysis was performed on 50,000 images (4 cancer types * 250 WSIs * 50 patches/WSI), split into 5 batches of 10,000 images each (50 WSIs/batch). The sample size and batching strategy was chosen because of the high computational costs of RSA, which grows according to $N^{2}$, and to provide a measure of uncertainly as quantified by the range across batches.
@@ 3.4 Representational Dissimilarity Matrix Generation
Representational dissimilarity matrix (RDM) construction consists of computing pairwise distances between groups of embeddings. We computed an RDM separately for each model using the Euclidean distance as the distance metric between embedding pairs. The Euclidean distance was used as it is recommended over correlation-based distances when computing RDMs given increased robustness to noise at the representation level (Botero2024; Bosch2025). When plotting the RDMs, the rows and columns were ordered by WSI and disease type (i.e., all patches from a given WSI are grouped together). RDMs were computed separately for each of the 5 data batches. The rsatoolbox Python package (10.1371/journal.pcbi.1003553) was used to compute the RDMs.
@@ 3.5 Comparing RDMs across Models
To quantify the similarity between model RDMs, we computed pairwise comparisons between all seven models. For each model pair, we loaded the upper-triangular values (excluding the diagonal) of their RDMs and computed the Spearman rank correlation. This process was performed separately for each of the 5 batches, yielding five similarity values per each model pair. The results are reported as the mean and range across these values. As an aggregate measure for each model, we also computed the average similarity to all other models. We note that the Spearman correlation is a common metric for comparing RDMs, preferred over Pearson correlation because it does not assume linearity (Bosch2025).
A key benefit of RSA is that it abstracts away the differences in embedding dimensionality between models. Specifically, as RSA uses RDMs for model comparison by quantifying pairwise similarities between embeddings across all input images, each RDM has dimensions $N\times N$, where $N$ is the size of the image set, independent of the original embedding size ($d$). These RDMs can then be directly compared across models (e.g. using Spearman correlation) to assess representational similarity. Consequently, RSA enables direct comparison of models with different embedding dimensionalities without requiring dimensional alignment or projection.
To quantitatively group different models according to similarity, we performed hierarchical clustering on the RDM similarity matrices. Ward’s method (Ward01031963) was used to produce a linkage tree, visualized as a dendrogram.
@@ 3.6 Assessing Disease and Slide Specificity
In more detail, for slide specificity, we computed the distribution of intra-slide distances as pairwise Euclidean distances between all 50 patch embeddings for a given slide, and aggregating across all diseases and slides. Inter-slide distances were computed by measuring Euclidean distances between all combinations of patches between different slides. Cliff’s Delta was computed between the intra- and inter-slide distance distributions, and this process was performed separately for each of the 5 batches. For each model, we reported the mean and range of Cliff’s Delta values across batches as an indicator of slide specificity.
@@ 4.1 Representational Similarity
By converting the representations for each model in a standardized format that is agnostic to feature size, the RDMs enable comparisons of representations across models. Figure 2 summarizes the degree of similarity between each pair of models, quantified as the Spearman correlation between the RDMs of each model. Higher similarities indicate that image patches with embeddings that are farther apart in one model also tend to be farther apart in the representational space of the other model, indicating common representational structures. The highest overall similarity is observed between UNI2 and Prov-Gigapath, with a Spearman correlation of 0.555 between their RDMs (range of 0.531-0.575 across 5 independent batches of $10K$ image patches). This finding is intuitive as UNI2 and Prov-Gigapath are both vision-only models trained using the Dinov2 algorithm. However, not all patterns align with similarities between self-supervised learning strategy. For instance, out of all of the foundation models, Virchow2’s lowest similarity was observed for UNI2, despite both using DinoV2 for training. On average, Virchow2 and UNI2 were the most distinct in their representational structure, with a mean correlation with the other CPath models of 0.419 and 0.421, respectively (Table 2). Prov-GigaPath had the highest average similarity (0.523). These patterns were consistent across each of the five independent data batches used in this study (each containing 10,000 images), where Virchow2 and UNI2 exhibited the lowest average correlations and Prov-Gigapath exhibited the highest correlation within each batch. Paired t-tests across batches confirmed that these differences were statistically significant for all comparisons ($p<0.05$ for all; i.e., Virchow2 << {Prov-GigaPath, CONCH, PLIP, KEEP}, UNI2 << {Prov-GigaPath, CONCH, PLIP, KEEP}, and Prov-GigaPath >> {UNI2, Virchow2, CONCH, PLIP, KEEP}). The baseline ViT-Dinov2 trained on natural images had low similarity ($\leq$0.3) to each model except PLIP (0.503). Hierarchical clustering on the Spearman similarity matrix revealed similar patterns (Figure 3), with ViT-Dinov2 and PLIP separating from the other foundation models.
@@ 4.2 Disease- and Slide- Specificity
In addition to model similarity, we evaluated the degree of slide- and disease-specificity for each model by comparing the relative distances between embeddings from the same slide and different slides, and separately from the same disease and different diseases. Cliff’s Delta, a non-parametric measure of the magnitude of difference between two groups, was used to quantify these effects. Ranging from 0 to 1, a Cliff’s Delta greater than 0.43 is regarded as a large effect, with values of 0.28 and 0.11 indicating medium and weak effects, respectively (vargha).
@@
The results of the slide- and disease-specificity analyses are contained in Table 3. All of the CPath models demonstrate relatively large slide-specificity, indicating that embeddings from patches from the same WSI tend to be closer together than embeddings from different WSIs. The highest level of slide specificity was observed for Prov-Gigapath, with a Cliff’s Delta of 0.762 (range: 0.658-0.807 across the 5 batches), followed closely by UNI2 (0.751, [0.685, 0.780]). CONCH demonstrated the lowest value of 0.596 [0.501, 0.636]. All of the CPath models had a higher slide-specificity than the ViT-Dinov2 baseline (0.367). For disease-specificity, the associations were weak for each of the models, with a range of 0.120 (Virchow) to 0.210 (KEEP). Thus, while embeddings from the same disease tended to be slightly closer together than embeddings from different diseases, the representations had a much stronger dependency on individual slides. No clear trends were observed regarding vision-language versus vision-only models for either slide- or disease-specificity.